# CMaP-SAM: Contraction Mapping Prior for SAM-driven Few-shot Segmentation

### v1

@@ Abstract
Few-shot segmentation (FSS) aims to segment new classes using few annotated images. While recent FSS methods have shown considerable improvements by leveraging Segment Anything Model (SAM), they face two critical limitations: insufficient utilization of structural correlations in query images, and significant information loss when converting continuous position priors to discrete point prompts. To address these challenges, we propose CMaP-SAM, a novel framework that introduces contraction mapping theory to optimize position priors for SAM-driven few-shot segmentation. CMaP-SAM consists of three key components: (1) a contraction mapping module that formulates position prior optimization as a Banach contraction mapping with convergence guarantees. This module iteratively refines position priors through pixel-wise structural similarity, generating a converged prior that preserves both semantic guidance from reference images and structural correlations in query images; (2) an adaptive distribution alignment module bridging continuous priors with SAM’s binary mask prompt encoder; and (3) a foreground-background decoupled refinement architecture producing accurate final segmentation masks. Extensive experiments demonstrate CMaP-SAM’s effectiveness, achieving state-of-the-art performance with 71.1 mIoU on PASCAL-$5^{i}$ and 56.1 on COCO-$20^{i}$ datasets.
@@
[label1]organization=University of Electronic Science and Technology of China, city=Chengdu, postcode=611731, country=China
@@ 1 Introduction
Few-shot segmentation (FSS) [5, 6, 7, 8, 9, 10] addresses these limitations by enabling models to segment novel object categories using only a handful of annotated examples. The core idea involves leveraging limited labeled samples in a support set to guide the segmentation of query images. In many FSS methods, position prior [5, 6] play an important role in directing segmentation. These priors, typically generated through feature matching or prototype learning between support and query images, provide probabilistic spatial representations that indicate potential target regions. By incorporating these position priors, models can more effectively filter out background distractions and focus on regions likely to contain the target object.
@@ 2.2 SAM-based Segmentation
The Segment Anything Model (SAM) [11] revolutionized zero-shot segmentation through its category-agnostic approach. Subsequent research advanced SAM’s capabilities along three strategy: semantic enhancement, domain specialization, and few-shot integration.Semantic enhancement addressed SAM’s limited categorical understanding through models like Semantic-SAM [28], which improved granularity via joint training, and OV-SAM [29], which incorporated CLIP-based [30] text prompting. Domain specialization tailored SAM to specific fields: RSPrompt [31] improved segmentation quality in remote sensing, MAS-SAM [32] was designed for marine animal segmentation, while SAM-Med2D [33], and SurgicalSAM [34] optimized performance for various medical imaging applications through specialized prompt encoders. Most recently, few-shot integration transformed SAM from requiring labor-intensive manual prompts to achieving automated segmentation with minimal examples. PerSAM [35] enabled single-sample model personalization, Matcher [36] developed robust bidirectional prompt sampling, and FSS-SAM [13] combined few-shot predictions with SAM processing.
@@ 3.3.1 Initial Prior Construction
[…]
@@ 3.3.2 Structure Transfer Matrix Construction
[…]
@@ 3.5 Foreground-Background Decoupled Refinement Module
[…]
@@ 4.1 Datasets and Evaluation Metrics
4.2 Implementation Details
Implementation Details: All experiments were implemented on the PyTorch framework  [^1] https://pytorch.org using 4 NVIDIA RTX 3090 GPUs. Following previous method [13], DINOv2 (ViT-Base-P14) [46] was employed as the feature extractor, with input images resized to 896$\times$896 pixels to extract 64$\times$64 feature maps. The SAM model operated at 1024$\times$1024 resolution, producing matching 64$\times$64 feature maps to maintain spatial consistency. Position prior generation was configured with 8 neighbors and a temperature coefficient of 0.1. Training was conducted using AdamW optimizer with a 1e-4 learning rate and batch size of 2 for 20 epochs on PASCAL-5i and FSS-1000, and 5 epochs on COCO-20i. Data augmentation included random horizontal flipping, scale transformations (0.8-1.2$\times$), rotation ($\pm$10°), and brightness/contrast adjustments. For K-shot evaluations, predictions were averaged as $\mathbf{M}_{\text{final}}=\frac{1}{K}\sum_{k=1}^{K}\mathbf{M}_{k}$.
@@ 4.3 Comparison with State-of-the-Art Methods
[…]
@@
[…]
@@
[…]
@@
[…]
@@ 4.4 Ablation Studies
Table 7: Sensitivity analysis with respect to the number of reference images $K$.
Quantity $\mathbf{K}$ mIoU $\triangle$ FB-IoU $\triangle$
1-shot 78.1 - 88.2 -
2-shot 80.3 2.2 89.8 1.6
3-shot 81.4 3.3 90.4 2.2
4-shot 82.3 4.2 90.9 2.7
5-shot 82.2 4.1 90.9 2.7
6-shot 82.2 4.1 90.9 2.7
7-shot 82.5 4.4 91.1 2.9
8-shot 82.4 4.3 91.1 2.9
9-shot 82.5 4.4 91.1 2.9
10-shot 82.5 4.4 91.1 2.9
@@ 4.6 Visualizations
Visualization of Segmentation Results: Fig. 5 illustrates segmentation results on the PASCAL-$5^{i}$ and COCO-$20^{i}$ datasets. The method exhibited precise capture of complex aircraft contours (row 1), effective segmentation of grayscale horse images despite limited color information (row 2), and robust handling of scale variations when segmenting dense sheep herds from a single reference instance (row 3). Notably, in the display screen example (row 4), the method produced semantically superior results compared to the ground truth, correctly excluding erroneously labeled hand regions. These visualizations validated the effectiveness of the contraction mapping theory-based position prior optimization framework across diverse and challenging scenarios.

### v2

@@ Abstract
Few-shot segmentation (FSS) aims to segment new classes using few annotated images. While recent FSS methods have shown considerable improvements by leveraging Segment Anything Model (SAM), they face two critical limitations: insufficient utilization of structural correlations in query images, and significant information loss when converting continuous position priors to discrete point prompts. To address these challenges, we propose CMaP-SAM, a novel framework that introduces contraction mapping theory to optimize position priors for SAM-driven few-shot segmentation. CMaP-SAM consists of three key components: (1) a contraction mapping module that formulates position prior optimization as a Banach contraction mapping with convergence guarantees. This module iteratively refines position priors through pixel-wise structural similarity, generating a converged prior that preserves both semantic guidance from reference images and structural correlations in query images; (2) an adaptive distribution alignment module bridging continuous priors with SAM’s binary mask prompt encoder; and (3) a foreground-background decoupled refinement architecture producing accurate final segmentation masks. Extensive experiments demonstrate CMaP-SAM’s effectiveness, achieving state-of-the-art performance with 71.1 mIoU on PASCAL-$5^{i}$ and 56.1 on COCO-$20^{i}$ datasets. Code is available at https://github.com/Chenfan0206/CMaP-SAM.
@@
[…]
@@ 1 Introduction
Few-shot segmentation (FSS) [5, 6, 7, 8, 9, 10, 11, 12, 13, 14] addresses these limitations by enabling models to segment novel object categories using only a handful of annotated examples. The core idea involves leveraging limited labeled samples in a support set to guide the segmentation of query images. In many FSS methods, position prior [5, 6] play an important role in directing segmentation. These priors, typically generated through feature matching or prototype learning between support and query images, provide probabilistic spatial representations that indicate potential target regions. By incorporating these position priors, models can more effectively filter out background distractions and focus on regions likely to contain the target object.
@@ 2.2 SAM-based Segmentation
The Segment Anything Model (SAM) [15] revolutionized zero-shot segmentation through its category-agnostic approach. Subsequent research advanced SAM’s capabilities along three strategy: semantic enhancement, domain specialization, and few-shot integration. Semantic enhancement addressed SAM’s limited categorical understanding through models like Semantic-SAM [32], which improved granularity via joint training, and OV-SAM [33], which incorporated CLIP-based [34] text prompting. Domain specialization tailored SAM to specific fields: RSPrompt [35] improved segmentation quality in remote sensing, MAS-SAM [36] was designed for marine animal segmentation, while SAM-Med2D [37], SurgicalSAM [38] optimized performance for various medical imaging applications through specialized prompt encoders. In addition, Hi-SAM [39] leveraged parameter-efficient fine-tuning for hierarchical text segmentation, and SAM-HQ [40] introduced a learnable high-quality output token to improve performance on high-resolution images. Most recently, few-shot integration transformed SAM from requiring labor-intensive manual prompts to achieving automated segmentation with minimal examples. PerSAM [41] enabled single-sample model personalization, Matcher [42] developed robust bidirectional prompt sampling, and FSS-SAM [17] combined few-shot predictions with SAM processing.
@@ 3.3.1 Initial Prior Construction
where $\text{sim}(\cdot,\cdot)$ denotes cosine similarity, $\langle\cdot,\cdot\rangle$ represents inner product, and $|\cdot|$ represents Euclidean norm. To enhance the discriminative power of the position prior, min-max normalization is applied to map the position prior to the $[0,1]$ range and amplifying the contrast between foreground and background regions:
@@ 3.3.2 Structure Transfer Matrix Construction
where $P_{i,j}$ represents elements in the similarity matrix, and $\text{top-}k(P_{i})$ denotes the set of $k$ elements with highest similarity values in row $i$. By setting non-Top-$k$ elements to negative infinity, their probability weights after softmax normalization approach zero, achieving strict sparsification. This design reduces computational complexity from $O(N_{q}^{2})$ to $O(kN_{q})$, significantly alleviating the computational burden when processing high-resolution feature maps.
@@ 3.5 Foreground-Background Decoupled Refinement Module
where $\|$ represents channel-dimension concatenation, and $\mathcal{M}(\cdot)$ is a multi-modal fusion function composed of gated attention units with weight coefficients $\alpha_{i}^{s}\in[0,1]$ calculated as:
@@ 4.1 Datasets and Evaluation Metrics
Evaluation Metrics: Performance assessment was carried out using cross-fold validation, where models were trained on three folds and tested on the remaining fold to evaluate their generalization capability to novel classes. Two complementary metrics were employed to ensure a comprehensive evaluation. The first metric, class-wise mean Intersection-over-Union (mIoU), calculated the average IoU across all test classes, defined as $\text{mIoU}=\frac{1}{C}\sum_{c=1}^{C}\text{IoU}_{c}$. The second metric, Foreground-Background IoU (FB-IoU), measured the average IoU between foreground and background regions, defined as $\text{FB-IoU}=\frac{1}{2}(\text{IoU}_{\text{F}}+\text{IoU}_{\text{B}})$.
4.2 Implementation Details
@@ 4.3 Comparison with State-of-the-Art Methods
Table 1: Comparison of 1-shot performance on $\text{PASCAL-}5^{i}$ dataset. Best in bold, second best in underlined.
@@
Table 2: Comparison of 5-shot performance on $\text{PASCAL-}5^{i}$ dataset. Best in bold, second best in underlined.
@@
Table 1 and Table 2 show the 1-shot and 5-shot performance on the $\text{PASCAL-}5^{i}$ dataset, respectively. The proposed CMaP-SAM method consistently demonstrated superior performance compared to existing state-of-the-art methods across various backbones. In the 1-shot setting, CMaP-SAM achieved an mIoU of 71.1 and an FB-IoU of 80.8, exceeding the previous best specialized method (ABCNet) by 5.5 points in mIoU. Additionally, CMaP-SAM outperformed the best general model (FSS-SAM) by 2.6 points in mIoU. In the 5-shot setting, CMaP-SAM further improved its performance, achieving an mIoU of 76.3 and an FB-IoU of 85.1.
Table 3: Comparison of 1-shot performance on $\text{COCO-}20^{i}$ dataset. Best in bold, second best in underlined.
@@
Table 4: Comparison of 5-shot performance on $\text{COCO-}20^{i}$ dataset. Best in bold, second best in underlined.
@@
Table 5 provides a detailed comparison of the proposed CMaP-SAM method with PGMA-Net [54] and GF-SAM [55] on both $\text{PASCAL-}5^{i}$ and $\text{COCO-}20^{i}$ datasets. The results indicate that CMaP-SAM outperforms PGMA-Net in both datasets, achieving an mIoU of 75.6 on $\text{PASCAL-}5^{i}$ and 60.6 on $\text{COCO-}20^{i}$ when using CLIP, while maintaining competitive performance without CLIP. Additionally, CMaP-SAM+ (using SAM-Huge) surpasses GF-SAM by 2.1 points on $\text{PASCAL-}5^{i}$ and 1.8 points on $\text{COCO-}20^{i}$, demonstrating the effectiveness of the proposed method in enhancing segmentation performance.
Table 5: In-depth comparison with PGMA-Net [54] and GF-SAM [55] on $\text{PASCAL-}5^{i}$ and $\text{COCO-}20^{i}$ datasets. Best in bold, second best in underlined.
Method Details $\text{PASCAL-}5^{i}$ $\text{COCO-}20^{i}$
mIoU mIoU
PGMA-Net [54] using CLIP 74.1 59.4
CMaP-SAM (Ours) no CLIP 71.1 56.1
CMaP-SAM* (Ours) using CLIP 75.6 60.6
GF-SAM [55] SAM-Huge 72.1 58.7
CMaP-SAM (Ours) SAM-Base 71.1 56.1
CMaP-SAM+ (Ours) SAM-Huge 74.2 60.5
@@ 4.4 Ablation Studies
Table 8: Sensitivity analysis with respect to the number of reference images $K$ on Pascal-5i (Fold-0) and COCO-20i (Fold-0).
Pascal-5i (Fold-0) COCO-20i (Fold-0)
Quantity $\mathbf{K}$ mIoU $\triangle$ FB-IoU $\triangle$ mIoU $\triangle$ FB-IoU $\triangle$
1-shot 78.1 88.2 54.0 73.5
2-shot 80.3 2.2 89.8 1.6 58.2 4.2 75.8 2.3
3-shot 81.4 3.3 90.4 2.2 59.3 5.3 76.9 3.4
4-shot 82.3 4.2 90.9 2.7 60.7 6.7 77.2 3.7
5-shot 82.2 4.1 90.9 2.7 61.3 7.3 77.5 4.0
6-shot 82.2 4.1 90.9 2.7 61.8 7.8 78.3 4.8
7-shot 82.5 4.4 91.1 2.9 61.8 7.8 78.3 4.8
8-shot 82.4 4.3 91.1 2.9 61.8 7.8 78.3 4.8
9-shot 82.5 4.4 91.1 2.9 61.8 7.8 78.3 4.8
10-shot 82.5 4.4 91.1 2.9 61.8 7.8 78.3 4.8
@@ 4.5 Sensitivity Analysis
To investigate whether the sensitivity to the number of support images varies across datasets, we extended the analysis to COCO-20i (Fold-0). The right part of Table 8 shows that the performance gain per additional shot is similar to that observed on PASCAL-5i, confirming the method’s cross-dataset insensitivity to support-set size.
Moreover, we performed another sensitivity analysis to evaluate the effect of stacking more convolution blocks on model performance. As shown in Table 9, increasing the number of blocks from 1 to 5 only yields marginal gains in mIoU (from 78.1% to 78.4%). This indicates that a single convolution block is sufficient for balancing performance and efficiency.
Table 9: Sensitivity analysis with respect to the complexity of the adapter (number of stacked convolution blocks inside the adapter).
Number of Blocks 1 2 3 4 5
Performance (mIoU) 78.1 78.1 78.3 78.4 78.4
4.6 Complexity analysis
Table 10 presents the computational complexity analysis, including the number of parameters, multiply-accumulate operations (MACs), and inference throughput (FPS). Notably, CMaP-SAM introduces only 2.0M trainable parameters, which is significantly fewer than competing methods such as DCAMA (14.2M). However, due to inheriting SAM’s large ViT-based image encoder, CMaP-SAM incurs higher total parameter count and MACs, resulting in a lower inference speed. This trade-off prioritizes segmentation accuracy over real-time efficiency. Therefore, CMaP-SAM is more suitable for offline processing scenarios, such as post-hoc analysis of medical imaging [37, 38], and the real-time deployment is left for future work via lightweight backbones (e.g., EfficientSAM [59]), pruning, or distillation.
Table 10: Complexity analysis of CMaP-SAM and comparison with other methods.
Method params (M) MACs Speed
learnable frozen total (G) (fps)
HSNet 2.6 25.6 28.2 20.1 16.0
DCAMA 14.2 25.6 39.8 39.8 19.7
PGMA-Net 2.7 23.6 26.3 42.3 10.5
CMaP-SAM(Ours) 2.0 177.2 179.2 87.6 1.9
@@ 4.7 Visualizations
Visualization of Segmentation Results: Fig. 5 illustrates segmentation results on the PASCAL-$5^{i}$ and COCO-$20^{i}$ datasets. The method demonstrated precise segmentation of object boundaries in simple cases (row 1), effective handling of sheep with significant scale variations and grayscale horse images despite limited color information (row 2), and robust performance in occlusion scenarios where occluded chair parts were accurately segmented while producing semantically superior results compared to the ground truth (e.g., correctly excluding erroneously labeled hand regions) (row 3). Notably, even in complex backgrounds, the method maintained accurate segmentation of chairs and cars (row 4). These visualizations validate the effectiveness of the contraction mapping theory-based position prior optimization framework across diverse and challenging scenarios. However, as shown in row 5, the method fails when the foreground and background objects in the query image exhibit similar appearances, leading to confusion. Furthermore, significant disparities between support and query images, such as those observed for plants with highly dissimilar appearances, adversely impact performance.
@@
4.8 Limitations and Future Work
Despite the promising performance of CMaP-SAM, several limitations remain. First, the model inherits the computational overhead from SAM, as it retains ViT-based image encoder, resulting in higher FLOPs compared to more lightweight CNN-based alternatives. Second, a performance saturation phenomenon is observed in our experiments: as the number of support samples increases, the improvement in performance gradually plateaus, which remains a critical challenge in FSS field. These limitations suggest directions for future work, including the exploration of lightweight SAM variants (e.g., EfficientSAM [59]), pruning, and model distillation techniques to reduce inference latency while maintaining high accuracy. Additionally, developing meta-learned prompt generators could further improve efficiency and scalability for real-world applications.
@@ 5 Conclusion
Acknowledgements
This work was supported in part by National Science and Technology Major Project (2021ZD0112001), National Natural Science Foundation of China (Grant 62271119), Natural Science Foundation of Sichuan Province (2023NSFSC1972).